New Stock Issue

Bynum and Crumpton Inc. (B&C), a small jewelry manufacturer, has been successful and has enjoyed a positive growth trend. Now B&C is planning to go public with an issue of common stock, and it faces the problem of setting an appropriate price for the stock. The company and its investment banks believe that the proper procedure is to conduct a valuation and select several similar firms with publicly traded common stock and to make relevant comparisons.
Several jewelry manufacturers are reasonably similar to B&C with respect to product mix, asset composition, and debt/equity proportions. Of these companies, Abercrombe Jewelers and Gunter Fashions are most similar. When analyzing the following data, assume that the most recent year has been reasonably "normal" in the sense that it was neither especially good nor especially bad in terms of sales, earnings, and free cash flows. Abercrombe is listed on the AMEX and Gunter on the NYSE, while B&C will be traded in the Nasdaq market.

Company data Abercrombe Gunter B&C
Shares outstanding 6 million 12 million 500,000
Price per share $30.00 $53.00 NA
Earnings per share $2.20 $3.13 $2.60
Free cash flow per share $1.63 $2.54 $1.90
Book value per share $15.00 $23.00 $19.00
Total assets $125 million $326 million $13.5 million
Total debt $35 million $50 million $4 million
  1. B&C is a closely held corporation with only 500,000 shares outstanding. Free cash flows have been low and in some years negative due to B&C's recent high sales growth rates, but as its expansion phase comes to an end B&C's free cash flows should increase. B&C anticipates the following free cash flows over the next 5 years:
    Year 1 2 3 4 5
    FCF $1,000,000 $1,050,000 $1,208,000 $1,329,000 $1,462,000

    After Year 5, free cash flow growth will be stable at 7% per year. Currently, B&C has no non-operating assets, and its WACC is 12%. Using the free cash flow valuation model, estimate B&C's intrinsic value of equity and intrinsic per share price. Do not round intermediate calculations. Round your answers for the value of equity to the nearest dollar and for the value of equity per share to the nearest cent.

    Value of equity

    $  

    Per share value of equity

    $  

  2. Calculate debt to total assets, P/E, market to book, P/FCF, and ROE for Abercrombe, Gunter, and B&C. For calculations that require a price for B&C, use the per share price you obtained with the corporate valuation model in Part a. Do not round intermediate calculations. Round your answers to two decimal places.

    Abercrombe Gunter B&C
    D/A % % %
    P/E
    Market/Book
    ROE % % %
    P/FCF
  3. Using Abercrombe's and Gunter's P/E, Market/Book, and Price/FCF ratios, calculate the range of prices for B&C's stock that would be consistent with these ratios. For example, if you multiply B&C's earnings per share by Abercrombe's P/E ratio you get a price. What range of prices do you get? Do not round intermediate calculations. Round your answers to the nearest cent.

    The range of prices:
    from $   to $  

    How does this compare with the price you get using the corporate valuation model?

    The price obtained with the corporate valuation model is -Select-withinout ofItem 20 this range of prices.
Answer #1

a) Value of equity:

Cash flow in year 5 = $1,462,000

given growth rate after year 5 = 7%

so cash flow from year 6 on wards = 1,462,000 + 7% = $1,564,340

using growth perpetuity formula = future cash flow / Ke - g

here,

future cash flow = 1,564,340

Ke = WACC = 12%

growth = g =7%

continuous value = 1564340 / 12% - 7%

= $31,286,800

calculating present values:

year FCF PV factor @12% PV of cash flows

1 1,000,000 0.89285714 892,857.143

2 1,050,000 0.797193877 837,053.571

3 1,208,000 0.711780247 859,830.539

4 1,329,000 0.635518078 844,603.526

5 1,462,000 0.5674268557 829,578.063

6 31,286,800    0.5674268557 17,752,970.549

present value 22,016,893.39

so from the above total value of B & C = $22,016,893(rounded off to nearest dollar)

value of debt = $4,000,000

so value of equity(total value - value of debt) = 22,016,893 - 4,000,000 = $18,016,893

number of shares outstanding = 500,000

so price per share = 18,016,893 / 500,000 = $36.034

b)

Abercrombe Gunter B & C
1)debt / assets (D / A) (35 / 125)*100 = 28% (50 / 326)*100 = 15.34% (4 / 13.5)*100 = 29.63%
2)P / E(market price / EPS) 30 / 2.2 = 13.64 53 / 3.13 = 16.93 36.034 / 2.6 = 13.86
3)MARKET / BOOK 30 / 15 = 2 53 / 23 = 2.30 36.034 / 19 = 1.90
4)ROE(EPS / BOOK VALUE) (2.2 / 15)*100 = 14.67% (3.13 / 23)*100 = 13.61% (2.6 / 19)*100 = 13.68
5)PRICE(Market) / FCF 30 / 1.63 = 18.40 53 / 2.54 = 20.87 36.034 / 1.90 = 18.97

C)

Range of prices:

using P/E ratio:

Abercrombe Gunter

P/E ratio 13.64 16.93

EPS Of B&C   2.6 2.6

Price of B&C 35.46 44.02

Price = P/E x EPS

Using M/B values:

Abercrombe Gunter

M/B ratio 2     2.3

Book value Of B&C 19 19

Price of B&C 38 43.7

Price = M/B ratio x book value of B&C

Using Price / FCF Ratios:

Abercrombe Gunter

Price/FCF ratio 18.40 20.87

FCF Of B&C 1.9 1.9   

Price of B&C 34.96 39.65

Using the above data

Range = lowest price to highest price

Range = 34.96 to 44.02

Price of B&C = $36.034

So the price is Within the range